# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

### v1

@@ 1.1 Our Contribution
$Y_{i}=\beta(A_{i})D_{i}+g(X_{i})+\epsilon_{i},\,i=1,\ldots,N$ (1)
for a response variable $Y_{i}$, a univariate continuous covariate $A_{i}$, the treatment variable $D_{i}$, covariates $X_{i}$ and an error term $\epsilon_{i}$. We focus here on univariate $A_{i}$, but the extension to the multivariate case is conceptually straightforward, see also the discussion in Section 5. The goal is to perform inference for $\beta(a)$ for some specific value $a$. The error $\epsilon_{i}$ satisfies $\mathbb{E}[\epsilon_{i}|X_{i}]=0$ but we have endogeneity, meaning that $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]\neq 0$. However, we assume that we have access to an instrumental variable $Z_{i}$ that satisfies $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$. Our method is based on double/debiased machine learning [12] and uses a kernel smoothing approach to model the heterogeneity in the continuous variable $A_{i}$. We give conditions under which our method yields consistency and asymptotically valid confidence sets. Our method is simple and the results are easy to interpret. Moreover, many extensions of our method are possible by replacing kernel smoothing with other machine learning methods. Along the way, we also provide a discussion of some aspects of double machine learning with instrumental variables for homogeneous treatment effects (i.e., $\beta(\cdot)\equiv\beta$ is a constant) which, to our knowledge, have not yet been comprehensively discussed. Concretely, we discuss the usage of machine learning instruments and weak-instrument-robust confidence intervals. Finally, we provide an implementation of our method in the R package IVDML, which is available on GitHub (https://github.com/cyrillsch/IVDML).
@@ 1.2 Partially Linear IV Models
If the treatment effect $\beta(\cdot)$ in (1) is constant (i.e., does not depend on $A_{i}$), the model reduces to a standard partially linear model with endogeneity (i.e. $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]\neq 0$)
@@ 1.3 Heterogeneous Treatment Effects
There is as well a large body of works, that consider (variants of) model (1) but assume no endogeneity, that is $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]=0$ and there is no need for an instrument. An overview can be found in Chapters 14 and 15 of [13]. Most works consider a binary treatment $D_{i}$ taking values in $\{0,1\}$. In that case, $\beta(A_{i})$ in (1), is the Conditional Average Treatment Effect (CATE) $\beta(A_{i})=\mathbb{E}[Y_{i}(1)-Y_{i}(0)|A_{i}]$, where $Y_{i}(1)$ and $Y_{i}(0)$ are the potential outcomes of $Y_{i}$ that would be observed when the treatment is assigned to be $D_{i}=0$ or $D_{i}=1$, respectively [13, 23]. We restrict ourselves here to the literature which is most related to our contribution.
A wide-spread approach for estimation and inference for the CATE is to express it as a conditional expectation $\beta(A_{i})=\mathbb{E}[W_{i}(\eta_{0})|A_{i}]$, where $W_{i}(\eta_{0})$ is a random variable depending on the observed quantities and the unknown potentially high-dimensional nuisance parameter $\eta_{0}$ which has to be estimated. If $W_{i}(\eta_{0})$ satisfies the Neyman orthogonality, one can use a DML approach with cross-fitting to obtain an estimate $W_{i}(\hat{\eta})$, which is then regressed on $A_{i}$ [13]. If $A_{i}$ is low-dimensional (as it is in our work), one can use standard methods and results for low-dimensional nonlinear regression to perform inference and estimation for the CATE. Such approaches are followed in [24] using series expansion, [17] using local linear regression, and [40] using locally constant regression. In [33], a modified version of the random forest algorithm is proposed, allowing to perform inference on heterogeneous treatment effects based on a similar idea. Another approach is to identify the heterogeneous treatment effect $\beta(A_{i})$ by a conditional moment restriction. This approach is particularly popular with random forest based methods, see [4, 22]. The above works mostly consider binary treatment $D$. Moreover, they do not cover the endogenous case, with the exception of [4, Sec.7], who apply their method to instrumental variables regression. Perhaps most related to our method is [28] which considers the endogenous setting and uses instrumental variables. The method also starts from a conditional moment equation similar to ours and then constructs a least squares objective which can be minimized by a user-chosen machine learning algorithm. Whereas the problem setup is the same, our method complements the work by [28] in several ways. To summarize, our approach is tailored to the special case of a treatment effect that depends smoothly on a univariate quantity, whereas their approach is more generic. With our method, we are able to perform inference for the treatment effect $\beta(a)$ at a concrete value $a$ and we extensively discuss the efficiency and both standard and weak-instrument robust inference. In contrast, [28] only briefly touch on inference for projections of the heterogeneous treatment effect on a simple space (e.g. the heterogeneity can be approximated by a linear function of a low number of covariates) and they do not explicitly discuss efficiency.
@@ 1.5 Notation
For two sequences $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ of positive real numbers, we write $a_{n}\lesssim b_{n}$ if there exists a constant $C>0$ such that for all $n\in\mathbb{N}$, $a_{n}\leq Cb_{n}$. We write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. We write $a_{n}\ll b_{n}$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. Moreover, we use stochastic order symbols: for a sequence of random variables $(X_{n})_{n\in\mathbb{N}}$ and a sequence of positive real numbers $(a_{n})_{n\in\mathbb{N}}$, we write $X_{n}=o_{P}(a_{n})$ if for all $\epsilon>0$, $\mathbb{P}(|X_{n}|/a_{n}\geq\epsilon)\to 0$, as $n\to\infty$. We write $X_{n}=O_{P}(a_{n})$ if $\limsup_{n\to\infty}\mathbb{P}(|X_{n}|/a_{n}\geq M)\to 0$, as $M\to\infty$.
@@ 2 Homogeneous Treatment Effect
Note that here, we do not have the restriction $d=1$ of univariate $Z$ anymore but the function $\zeta$ should be real-valued. A natural goal is to find a transformation $\zeta$ that is easy to obtain and such that the resulting estimator has a small variance. We use $\zeta(Z_{i},X_{i})=f(Z_{i},X_{i})\coloneqq\mathbb{E}[D_{i}|Z_{i},X_{i}]$, which can be easily obtained by nonlinear regression of $D_{i}$ versus $(Z_{i},X_{i})$ and coincides with the semiparametrically efficient instrument (see e.g. equation (9) in [31]) for the homoscedastic case $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathrm{const.}$ This means that under homoscedasticity, the estimator based on the identification (5) achieves minimum asymptotic variance with $\zeta=f$ (see also Section 2.5 below).
@@ 2.1 Construction
Fix $K\in\mathbb{N}$. Let $\left\{I_{k}\right\}_{k=1}^{K}$ be a $K$-fold random partition of the observation indices $[N]=\{1,\ldots,N\}$ such that each fold has size $|I_{k}|=n=N/K$ (for simplicity, we assume that $N/K$ is an integer). For $k=1,\ldots,K$, let $\hat{f}^{k}$, $\hat{\phi}^{k}$ and $\hat{l}^{k}$ be estimates of $f,\phi,l$ obtained using machine learning on the data $I_{k}^{c}$ (i.e. on all data apart from the $k$th fold). Define the residuals
@@
$R_{D,i}^{k}$ $\coloneqq D_{i}-\hat{\phi}^{k}(X_{i}),\,i\in I_{k},\,k=1,\ldots,K$ (7)
$R_{f,i}^{k}$ $\coloneqq\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}^{k}(X_{i}),\,i\in I_{k},\,k=1,\ldots,K.$ (8)
@@
Similarly to [2], we recommend not to use the direct ML estimator $\hat{\phi}$ in the definition of $R_{f,i}^{k}$ but instead use a two stage approach: for $k=1,\ldots,K$, obtain $\hat{f}^{k}$ as before by a nonlinear regression of $(D_{i})_{i\in I_{k}^{c}}$ versus $(Z_{i},X_{i})_{i\in I_{k}^{c}}$. In a second step, also regress $(\hat{f}^{k}(Z_{i},X_{i}))_{i\in I_{k}^{c}}$ on $(X_{i})_{i\in I_{k}^{c}}$ to obtain $\hat{\phi}^{k}_{2}$, which is a different estimate of $\phi$ than $\hat{\phi}^{k}$ and use
$R_{f,i}^{k}=\hat{f}(Z_{i},X_{i})-\hat{\phi}^{k}_{2}(X_{i}),\,i\in I_{k},\,k=1,\ldots,K.$ (10)
in the estimator (9). Note that we still use the direct estimate $\hat{\phi}^{k}$ in the definition (7) of $R_{D,i}^{k}$. The intuition behind using (10) instead of (8) is that due to misspecification it might happen that $\hat{f}^{k}$ does not converge to the true $f$ but to some function $\zeta$. Then, we are in the setting of the identification (5) and should subtract an estimate of $\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$ rather than an estimate of $\phi(X_{i})=\mathbb{E}[D_{i}|X_{i}]$. Hence, we expect the estimator $\hat{\beta}$ with $R_{f,i}^{k}$ based on (10) to be more robust against poor estimation of $f$.
@@ 2.3 Robust Confidence Intervals
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of a standard Gaussian distribution. Hence, the set $\mathcal{C}_{N}^{\alpha}=\left\{\beta:|\hat{Q}(\beta)|\leq z_{1-\alpha/2}\widehat{SE}_{Q}(\beta)/\sqrt{N}\right\}$ is an asymptotic $(1-\alpha)$-confidence set for $\beta$. The set $\mathcal{C}_{N}^{\alpha}$ is defined as the sublevel set of a parabola in $\beta$. Hence it is possible to derive an explicit expression, see Appendix C.1 for the expression in the heterogeneous case (the homogeneous case is analogue). Moreover, one can show that under some assumptions on the strength of the IV, the confidence interval $\mathcal{C}_{N}^{\alpha}$ is asymptotically equivalent to the classical confidence interval given in (12), see Appendix C.2 for the derivation in the heterogeneous case (the homogeneous case is analogue). Hence, one does not lose asymptotic efficiency from using the more robust version and we would recommend to use the robust confidence sets in practice. However, from a purely theoretical point of view, the advantage is not large: we still need that the nuisance function estimates converge fast enough, and essentially, the conditions on the convergence of the nuisance function estimates get scaled by the IV strength, which means that with weak IV, we get stronger requirements on how fast the nuisance function estimates need to converge, see also Remark 3.
@@ 2.4 Summary of Results for the Homogeneous Treatment Effect
which can be seen as a measure of IV strength on the population level. Note that $\sigma_{Q}^{2}$ is unnormalized, i.e. we do not divide by the noise level $\mathbb{E}[\epsilon_{i}^{2}]$. In contrast to the standard theory from [12] (Assumption 3.4. (d) there), we do not assume that $\sigma_{Q}^{2}$ is bounded away from 00, but we have to rescale the conditions on the convergence rates. We write $\|f-\hat{f}^{k}\|_{L_{2}}$ for the $L_{2}$-norm of $f-\hat{f}^{k}$. Note that $f$ is fixed and $\hat{f}^{k}$ depends on the data in $I_{k}^{c}$. Hence, $\|f-\hat{f}^{k}\|_{L_{2}}$ is random. More precisely, it is given by
@@
where $Z$ and $X$ are drawn from model (3) independently from $\{X_{i},Y_{i},Z_{i},D_{i}\}_{i\in I_{k}^{c}}$ and by conditioning on $I_{k}^{c}$ we mean conditioning on $\{X_{i},Y_{i},Z_{i},D_{i}\}_{i\in I_{k}^{c}}$. The analogous reasoning applies to $\|\phi-\hat{\phi}^{k}\|_{L_{2}}$ and $\|l-\hat{l}^{k}\|_{L_{2}}$. The main requirements for the nuisance function estimators are that for $k=1,\ldots,K$, it holds that
$\frac{1}{\sigma_{Q}}\|l-\hat{l}^{k}\|_{L_{2}}\|\phi-\hat{\phi}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right),$ (16)
@@
$\frac{1}{\sigma_{Q}}\|\phi-\hat{\phi}^{k}\|_{L_{2}}^{2}$ $=o_{P}\left(N^{-1/2}\right),$
$\frac{1}{\sigma_{Q}}\|\phi-\hat{\phi}^{k}\|_{L_{2}}\|f-\hat{f}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right).$
Note that for simplicity, we assume that we do not use the two-step procedure according to Remark 2 but only use the direct estimate of $\phi$. If $\sigma_{Q}^{2}$ is bounded from below, equation (16) is satisfied if all the nuisance function estimators converge at a rate of $o_{P}(N^{-1/4})$ which is comparably slow and satisfied by many modern machine learning methods, when assuming appropriate sparsity and/or smoothness conditions (we refer to [12] for a list of references). If on the other hand, $\sigma_{Q}^{2}$ is decreasing with $N$, which corresponds to the setting of weak IV, the machine learners need to converge at faster rates.
@@
If the nuisance functions were known, the robust confidence set construction from Section 2.3, would yield asymptotically valid inference regardless of the strength of the instruments. However, this nice property does not carry over if the nuisance functions are estimated, since in (16), the convergence rates are also scaled by the strength of the IV. The fastest convergence rate for the nuisance function estimates that we can hope for is an $L_{2}$ error of $O_{P}(N^{-1/2})$ (linear regression). Hence, in the best possible case, it still follows from (16) that $\sigma_{Q}^{2}\gg 1/N$. If we additionally assume that $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]$ is bounded away from 00 and $\infty$ independently of $N$, it follows that $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}]\asymp\sigma_{Q}^{2}\gg 1/N$. This rules out the classical “weak IV asymptotics” regime [26] which would correspond to $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}]\asymp 1/N$. In conclusion, from a theoretical point of view, one does not gain very much from using the robust confidence set construction compared to the standard confidence interval, which is in contrast to the classical treatment of weak IV. The reason for this is that compared to the classical treatment, we need to estimate nuisance functions.

### v2

@@ 1.1 Our Contribution
$Y_{i}=\beta(V_{i})D_{i}+g(X_{i})+\epsilon_{i},\,i=1,\ldots,N$ (1)
for a response variable $Y_{i}$, a univariate continuous covariate $V_{i}$, the treatment variable $D_{i}$, covariates $X_{i}$ and an error term $\epsilon_{i}$. We focus here on univariate $V_{i}$, but the extension to the multivariate case is conceptually straightforward, see also the discussion in Section 5. The goal is to perform inference for $\beta(v)$ for some specific value $v$. The error $\epsilon_{i}$ satisfies $\mathbb{E}[\epsilon_{i}|X_{i}]=0$ but we have endogeneity, meaning that $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]\neq 0$. However, we assume that we have access to an instrumental variable $Z_{i}$ that satisfies $\mathbb{E}[\epsilon_{i}|Z_{i},X_{i}]=0$. Our method is based on double/debiased machine learning [16] and uses a kernel smoothing approach to model the heterogeneity in the continuous variable $V_{i}$. We give conditions under which our method yields consistency and asymptotically valid confidence sets. Our main focus is neither to provide generic meta-algorithms for heterogenous treatment effect estimation in the spirit of [33, 25, 44] nor a detailed investigation of optimality properties [29, 30]. We rather want to introduce a simple, interpretable and easy-to-use method that works with user-chosen machine learning for nuisance function estimation. Nevertheless, many extensions of our method are possible by replacing kernel smoothing with other nonparametric regression or machine learning methods. Along the way, we also provide a discussion of some aspects of double machine learning with instrumental variables for homogeneous treatment effects (i.e., $\beta(\cdot)\equiv\beta$ is a constant) which, to our knowledge, have not yet been comprehensively discussed. Concretely, we discuss the usage of machine learning instruments and weak-instrument-robust confidence intervals. Finally, we provide an implementation of our method in the R package IVDML, which is available on CRAN [38].
@@ 1.2 Partially Linear IV Models
If the treatment effect $\beta(\cdot)$ in (1) is constant (i.e., does not depend on $V_{i}$), the model reduces to a standard partially linear model with endogeneity (i.e. $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]\neq 0$)
@@
1.3 Varying Coefficient Models
Another type of models related to (1) are varying coefficient models, see, e.g., [18, 27, 22, 23]. These are linear models, where the coefficient vector can vary depending on an additional variable, i.e.,
$Y_{i}=\gamma(U_{i})^{T}W_{i}+\epsilon_{i}.$ (3)
Moreover, varying coefficient models have also been studied in the endogenous setting with instrumental variables [9, 8, 43]. Structurally, our model (3) is related to model (1) by setting $U_{i}=(V_{i},X_{i})$, $W_{i}=(D_{i},1)$ and $\gamma(\cdot)=(\beta(\cdot),g(\cdot))$. However, there are important differences: we are only interested in conducting inference for the single component function $\beta(\cdot)$, whereas $g(\cdot)$ is a (potentially high-dimensional) nuisance function that allows to include additional control variables in a flexible fashion with arbitrary machine learning. We are not interested in estimating or conducting inference for the function $g(\cdot)$. In contrast, with varying coefficient models, typically, all the component functions are of equal interest and are modeled using low-dimensional techniques. Hence, we will look at model (1) not through the lens of varying coefficient models but heterogeneous treatment effects instead.
@@ 1.4 Heterogeneous Treatment Effects
When the heterogeneous treatment effect $\beta(\cdot)$ is of interest, there is a large body of works, that consider (variants of) model (1) but assume no endogeneity, that is $\mathbb{E}[\epsilon_{i}|D_{i},X_{i}]=0$ and there is no need for an instrument. An overview can be found in Chapters 14 and 15 of [17]. Most works consider a binary treatment $D_{i}$ taking values in $\{0,1\}$. In that case, $\beta(V_{i})$ in (1), is the Conditional Average Treatment Effect (CATE) $\beta(V_{i})=\mathbb{E}[Y_{i}(1)-Y_{i}(0)|V_{i}]$, where $Y_{i}(1)$ and $Y_{i}(0)$ are the potential outcomes of $Y_{i}$ that would be observed when the treatment is assigned to be $D_{i}=0$ or $D_{i}=1$, respectively [17, 37]. We restrict ourselves here to the literature which is most related to our contribution.
[…]
@@ 1.6 Notation
For two sequences $(a_{n})_{n\in\mathbb{N}}$ and $(b_{n})_{n\in\mathbb{N}}$ of positive real numbers, we write $a_{n}\lesssim b_{n}$ if there exists a constant $C>0$ such that for all $n\in\mathbb{N}$, $a_{n}\leq Cb_{n}$. We write $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$. We write $a_{n}\ll b_{n}$ if $a_{n}/b_{n}\to 0$ as $n\to\infty$. Moreover, we use stochastic order symbols: for a sequence of random variables $(X_{n})_{n\in\mathbb{N}}$ and a sequence of positive real numbers $(a_{n})_{n\in\mathbb{N}}$, we write $X_{n}=o_{P}(a_{n})$ if for all $\epsilon>0$, $\mathbb{P}(|X_{n}|/a_{n}\geq\epsilon)\to 0$, as $n\to\infty$. We write $X_{n}=O_{P}(a_{n})$ if $\limsup_{n\to\infty}\mathbb{P}(|X_{n}|/a_{n}\geq M)\to 0$, as $M\to\infty$. Equalities involving conditional expectations should be understood to hold almost surely.
@@ 2 Homogeneous Treatment Effect
Note that here, we do not have the restriction $d=1$ of univariate $Z_{i}$ anymore but the function $\zeta$ should be real-valued. A natural goal is to find a transformation $\zeta$ that is easy to obtain and such that the resulting estimator has a small variance. We use $\zeta(Z_{i},X_{i})=f(Z_{i},X_{i})\coloneqq\mathbb{E}[D_{i}|Z_{i},X_{i}]$, which can be easily obtained by nonlinear regression of $D_{i}$ versus $(Z_{i},X_{i})$ and coincides with the semiparametrically efficient instrument (see e.g. equation (9) in [47]) for the homoscedastic case $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]=\mathrm{const.}$ This means that under homoscedasticity, the estimator based on the identification (6) achieves minimum asymptotic variance with $\zeta=f$ (see also Section 2.5 below).
@@ 2.1 Construction
Fix $K\in\mathbb{N}$ (independent of $N$). Let $\left\{I_{k}\right\}_{k=1}^{K}$ be a $K$-fold random partition of the observation indices $N]=\{1,\ldots,N\}$ such that each fold has size $|I_{k}|=n=N/K$ (for simplicity, we assume that $N/K$ is an integer). For $k=1,\ldots,K$, let $\hat{f}^{k}$ and $\hat{l}^{k}$ be estimates of $f$ and $l$ and let $\hat{\phi}_{1}^{k}$ and $\hat{\phi}_{2}^{k}$ be estimates of $\phi$, where $\hat{f}^{k}$, $\hat{l}^{k}$, $\hat{\phi}_{1}^{k}$ and $\hat{\phi}_{2}^{k}$ are obtained by machine learning only using the data $I_{k}^{c}$ (i.e. on all data apart from the $k$th fold). The reason for having two different estimators $\hat{\phi}_{1}^{k}$ and $\hat{\phi}_{2}^{k}$ for $\phi$ will be made clear below in Remark 2.
Define the residuals
@@
$R_{D,i}^{k}$ $\coloneqq D_{i}-\hat{\phi}_{1}^{k}(X_{i}),\,i\in I_{k},\,k=1,\ldots,K$ (8)
$R_{f,i}^{k}$ $\coloneqq\hat{f}^{k}(Z_{i},X_{i})-\hat{\phi}_{2}^{k}(X_{i}),\,i\in I_{k},\,k=1,\ldots,K.$ (9)
@@
The estimators $\hat{f}^{k}$, $\hat{l}^{k}$ and $\hat{\phi}_{1}^{k}$ can be obtained by directly fitting a machine learning model of $D_{i}$ vs. $(Z_{i},X_{i})$, $Y_{i}$ vs. $X_{i}$ and $D_{i}$ vs. $X_{i}$, respectively, using only $i\in I_{k}^{c}$. In principle, one could use the same direct estimator $\hat{\phi}_{2}^{k}=\hat{\phi}_{1}^{k}$ in the definition of the residuals $R_{f,i}^{k}$. However, similarly to [2], we recommend to use a two-stage approach for $\hat{\phi}_{2}^{k}$: after obtaining $\hat{f}^{k}$ on $I_{k}^{c}$, regress $(\hat{f}^{k}(Z_{i},X_{i}))_{i\in I_{k}^{c}}$ on $(X_{i})_{i\in I_{k}^{c}}$ to obtain $\hat{\phi}^{k}_{2}$, which is a different estimate of $\phi$ than $\hat{\phi}_{1}^{k}$. The intuition behind using $\hat{\phi}_{k}^{2}$ instead of $\hat{\phi}_{k}^{1}$ in the definition (9) of $R_{f,i}^{k}$ is that due to misspecification it might happen that $\hat{f}^{k}$ does not converge to the true $f$ but to some function $\zeta$. Then, we are in the setting of the identification (6) and should subtract an estimate of $\mathbb{E}[\zeta(Z_{i},X_{i})|X_{i}]$ rather than an estimate of $\phi(X_{i})=\mathbb{E}[D_{i}|X_{i}]$. Hence, we expect the estimator $\hat{\beta}$ with $R_{f,i}^{k}$ based on $\hat{\phi}_{2}^{k}$ to be more robust against poor estimation of $f$.
@@ 2.3 Robust Confidence Intervals
where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of a standard Gaussian distribution. Hence, the set $\mathcal{C}_{N}^{\alpha}=\left\{\beta:|\hat{Q}(\beta)|\leq z_{1-\alpha/2}\widehat{SE}_{Q}(\beta)/\sqrt{N}\right\}$ is an asymptotic $(1-\alpha)$-confidence set for $\beta$. The set $\mathcal{C}_{N}^{\alpha}$ is defined as the sublevel set of a parabola in $\beta$. Hence it is possible to derive an explicit expression, see Appendix C.1 for the expression in the heterogeneous case (the homogeneous case is analogue). Moreover, one can show that under some assumptions on the strength of the IV, the confidence interval $\mathcal{C}_{N}^{\alpha}$ is asymptotically equivalent to the classical confidence interval given in (12), see Appendix C.2 for the derivation in the heterogeneous case (the homogeneous case is analogue). Hence, one does not lose asymptotic efficiency from using the more robust version and we would recommend to use the robust confidence sets in practice. However, from a purely theoretical point of view, the advantage is not large: we still need that the nuisance function estimates converge fast enough, and essentially, the conditions on the convergence of the nuisance function estimates get scaled by the IV strength, which means that with weak IV, we get stronger requirements on how fast the nuisance function estimates need to converge, see also Remark 3. Nevertheless, in simulations one can observe that the robust confidence intervals do not suffer from undercoverage in settings with weak IV where the coverage of the standard confidence intervals breaks down (see Appendix E.1).
@@ 2.4 Summary of Results for the Homogeneous Treatment Effect
which can be seen as a measure of IV strength on the population level. Note that $\sigma_{Q}^{2}$ is unnormalized, i.e. we do not divide by the noise level $\mathbb{E}[\epsilon_{i}^{2}]$. In contrast to the standard theory from [16] (Assumption 3.4. (d) there), we do not assume that $\sigma_{Q}^{2}$ is bounded away from 0, but we have to rescale the conditions on the convergence rates. We write $\|f-\hat{f}^{k}\|_{L_{2}}$ for the $L_{2}$-norm of $f-\hat{f}^{k}$. Note that $f$ is fixed and $\hat{f}^{k}$ depends on the data in $I_{k}^{c}$. Hence, $\|f-\hat{f}^{k}\|_{L_{2}}$ is random. More precisely, it is given by
@@
where $Z$ and $X$ are drawn from model (4) independently from $\{X_{i},Y_{i},Z_{i},D_{i}\}_{i\in I_{k}^{c}}$ and by conditioning on $I_{k}^{c}$ we mean conditioning on $\{X_{i},Y_{i},Z_{i},D_{i}\}_{i\in I_{k}^{c}}$. The analogous reasoning applies to $\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2}}$, $\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}$ and $\|l-\hat{l}^{k}\|_{L_{2}}$. The main requirements for the nuisance function estimators are that for $k=1,\ldots,K$, it holds that
$\frac{1}{\sigma_{Q}}\|l-\hat{l}^{k}\|_{L_{2}}\|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right),$ (16)
@@
$\frac{1}{\sigma_{Q}}\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2}}|\phi-\hat{\phi}_{2}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right),$
$\frac{1}{\sigma_{Q}}\|\phi-\hat{\phi}_{1}^{k}\|_{L_{2}}\|f-\hat{f}^{k}\|_{L_{2}}$ $=o_{P}\left(N^{-1/2}\right).$
If $\sigma_{Q}^{2}$ is bounded from below, equation (16) is satisfied if all the nuisance function estimators converge at a rate of $o_{P}(N^{-1/4})$ which is comparably slow and satisfied by many modern machine learning methods, when assuming appropriate sparsity and/or smoothness conditions. If on the other hand, $\sigma_{Q}^{2}$ is decreasing with $N$, which corresponds to the setting of weak IV, the machine learners need to converge at faster rates, see also Remark 3 below.
@@
If the nuisance functions were known, the robust confidence set construction from Section 2.3, would yield asymptotically valid inference regardless of the strength of the instruments. However, this nice property does not carry over if the nuisance functions are estimated, since in (16), the convergence rates are also scaled by the strength of the IV. The fastest convergence rate for the nuisance function estimates that we can hope for is an $L_{2}$ error of $O_{P}(N^{-1/2})$ (linear regression). Hence, in the best possible case, it still follows from (16) that $\sigma_{Q}^{2}\gg 1/N$. If we additionally assume that $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]$ is bounded away from 0 and $\infty$ independently of $N$, it follows that $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}]\asymp\sigma_{Q}^{2}\gg 1/N$. This rules out the classical “weak IV asymptotics” regime [41] which would correspond to $\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}]\asymp 1/N$. The actual amount of weakness (in terms of the order of $\sigma_{Q}^{2}$) that one can in theory allow for, depends on what rates the machine learning algorithm can achieve. For example, for sparse high-dimensional linear models, one can obtain an $L_{2}$ error of $O_{P}(\sqrt{\log p/N})$ [7], so that the requirement on the IV strength is $\sigma_{Q}^{2}\gg(\log p)^{2}/N$, which is only slightly weaker than classical weak IV asymptotics. In contrast, with more flexible models such as (nonlinear) sparse high-dimensional additive models [57], neural networks [14] or (variants of) random forest [50], one can afford less weakness. In conclusion, from the analysis based on theoretical bounds one does not gain very much from using the robust confidence set construction compared to the standard confidence interval, which is in contrast to the classical treatment of weak IV. The reason for this is that compared to the classical theory, we need to estimate nuisance functions. However, for finite sample sizes and in settings with weak instruments, the robust confidence sets exhibit superior empirical coverage compared to the standard confidence intervals in simulations (see Appendix E.1).